# Meson spectroscopy in the $Sp(4)$ gauge theory with three antisymmetric fermions

### v1

@@ II.1 Lattice action
$D^{(as)}\Psi_{j}(x)\equiv(4/a+m_{0})\Psi_{j}(x)-\frac{1}{2a}\sum_{\mu}\left\{\frac{}{}(1-\gamma_{\mu})U^{\rm{(as)}}_{\mu}(x)\Psi_{j}(x+\hat{\mu})+(1+\gamma_{\mu})U^{\rm{(as)}}_{\mu}(x-\hat{\mu})\Psi_{j}(x-\hat{\mu})\frac{}{}\right\}\,.$ (7)
@@ IV.2 Two-point correlation functions
We first consider $Z_{2}\otimes Z_{2}$ stochastic wall sources, with the one-end trick—see, e.g., Ref. [289]—implemented in the HiRep code777 Specifically, commit ID 1b204b6 [366]. . In comparison with using naive point sources (delta functions), this choice allows to improve the signal, and determine masses and decay constants with higher precision. We set the number of hits to $3$, for all meson correlation functions and for all the ensembles listed in Table II.1. We verified that this strategy leads to a reduction of statistical noise and the appearance of a cleaner plateau in the effective mass plots at large Euclidean time, compared to the point source, but for the same computational cost.
@@
Both APE and Wuppertal smearing have been recently implemented in the HiRep code, extensively tested and used for the measurements of spin-$1/2$ chimera baryons in quenched $Sp(4)$ gauge theories [153, 356]. Their use in the study of connected diagram contributions to flavour-singlet meson correlation functions in the $Sp(4)$ theory with $N_{\rm f}=2$ fundamental fermions dramatically improves noise control, and makes it possible to perform measurements in this sector of the theory [170]. Application to the $Sp(4)$ theory coupled to $N_{\rm f}=2$ and $N_{\rm as}=3$ fermions [156] provides a supporting tool for the spectral density approach to lattice spectroscopy [155].
@@ V.1 Masses and decay constants
We have evidence in our numerical results suggesting that in the limit of massless hyperquarks of the theory, confinement is taking place, with spontaneous breaking of the (approximate) non-Abelian global symmetry. [^8] As we have clear evidence of symmetry breaking, even in naive extrapolations to the massless limit, we do not discuss symmetric mass generation—see Ref. [370] and references therein. Yet, despite the wide range of hyperquark masses we explored, our results suggest that we expect to see significant deviations from the predictions of leading order (LO) chiral perturbation theory, as none of the ensembles are close enough to the massless limit. We devote the next paragraphs to presenting this evidence and discussing its implications.
@@
In Fig. 10 we combine our measurements of the masses of ${\rm ps}$, ${\rm v}$, ${\rm s}$, ${\rm av}$, and ${\rm at}$ mesons, and present them in units of the decay constant of the ${\rm ps}$ meson, $f_{\rm ps}$. We display the results as a function of $am_{\rm PCAC}$. The figure highlights the contrast between the scaling properties of the ${\rm ps}$ meson, and all the other mesons. This observation further enforces the conclusion that the massless, continuum extrapolation of this theory is not IR conformal, and hence justifies the adoption of an analysis strategy that assumes confinement and spontaneous symmetry breaking of (approximate) global symmetry.
@@ V.3 More about vector mesons
Even more speculatively, this ratio, ${m_{{\rm v}}}/{f_{{\rm ps}}}$, can be used to provide a rough estimate of the on-shell coupling constant associated with the decay of the vector into two pseudoscalar mesons, $g_{\rm v\,ps\,ps}$, via the phenomenological KSRF relation, $g_{\rm v\,ps\,ps}=m_{\rm v}/\sqrt{2}f_{\rm ps}$ [367, 368]. In the continuum and massless limits, our numerical results yield $g_{\rm v\,ps\,ps}=4.658(78)$, which is smaller than the QCD value of $\sim 5.9$ [375] and the one in the $Sp(4)$ gauge theory with two fundamental fermions, $m_{\rm V}/\sqrt{2}f_{\rm PS}=5.72(18)(13)$ [144]. It would be interesting to perform a direct measurement of the coupling $g_{\rm v\,ps\,ps}$, which requires to study the theory with lower hyperquark masses.
@@ VI Summary and Outlook
Table 10: Ratio of the masses of the flavored mesons included in the continuum and massless extrapolations, to the pseudoscalar decay constant, $m_{M}/f_{\rm ps}$. The masses are measured by applying APE and Wuppertal smearing in the definition of the correlation functions. The mass of the first excited state of vector mesons, ${\rm v}^{\prime}$, is determined through a GEVP analysis that includes the two, inequivalent, vector and tensor meson operators. We also report the ratio of the masses of the first-excited and ground state of the vector meson, $m_{{\rm v}^{\prime}}/m_{\rm v}$.
@@ Appendix A More numerical results
In this appendix, we summarise all the new numerical results on lattice measurements of meson masses and decay constants which have been carried out for this work. We adopted two different strategies for the analysis of the correlation function. Measurement extracted from correlation functions defined with $Z_{2}\otimes Z_{2}$ stochastic wall sources and sinks, for all the ensembles listed in Table II.1, are tabulated in Tables  VI and  VI, and are discussed in Sect. IV and Appendix C. We also report the results for the PCAC (unrenormalised) hyperquark mass, $am_{\rm PCAC}$, and for the quantities, $m_{\rm ps}L$ and $f_{\rm ps}L$. To facilitate comparison with phenomenological models and applications to model building, we find it useful to express the measurements also in units of the decay constant of the pseudoscalar meson, $f_{\rm ps}$, in Table VI.
As explained in the body of the paper, we select a subset of the ensembles, which we use for the massless and continuum extrapolation, and apply a combination of APE and Wuppertal smearing, to repeat the mass measurements, performing also the extraction of the first-excited states of vector mesons with the GEVP method. The results are tabulated in Table VI, and discussed in Secs. V.2 and V.3. We express them in units of $f_{\rm ps}$, in Table VI, in which we also report the ratio between the first excited and ground states of vector meson.
@@ Appendix D PCAC mass
We conclude by repeating that even if the dEFT fit to the data yields intriguing results, it remains a somewhat speculative exercise. Besides the aforementioned fact that we do not know whether a light scalar singlet is present in the spectrum, this analysis relies on the use of continuum relations, while we know that the values of $\beta$ available are such that significant corrections are expected to be for this analysis due to discretisation effects. Another concern might originate from the observation that the quality of the dEFT fit worsens as the $\beta$ value increases. Nevertheless, the results are encouraging, and it would be interesting to redo such analysis on lattice data that is closer to both the continuum and the massless limits.

### v2

@@ II.1 Lattice action
$D^{(as)}\Psi_{j}(x)\equiv(4/a+m_{0})\Psi_{j}(x)-\frac{1}{2a}\sum_{\mu}\left\{\frac{}{}(1-\gamma_{\mu})U^{\rm{(as)}}_{\mu}(x)\Psi_{j}(x+\hat{\mu})+(1+\gamma_{\mu})U^{\rm{(as)}\,\dagger}_{\mu}(x-\hat{\mu})\Psi_{j}(x-\hat{\mu})\frac{}{}\right\}\,.$ (7)
@@ IV.2 Two-point correlation functions
We first consider $Z_{2}\otimes Z_{2}$ stochastic wall sources, with the one-end trick—see, e.g., Ref. [289]—implemented in the HiRep code. In comparison with using naive point sources (delta functions), this choice allows to improve the signal, and determine masses and decay constants with higher precision. We set the number of hits to $3$, for all meson correlation functions and for all the ensembles listed in Table II.1. We verified that this strategy leads to a reduction of statistical noise and the appearance of a cleaner plateau in the effective mass plots at large Euclidean time, compared to the point source, but for the same computational cost.
@@
Both APE and Wuppertal smearing have been recently implemented in the HiRep code777 Specifically, commit ID 1b204b6 [366]. , extensively tested and used for the measurements of spin-$1/2$ chimera baryons in quenched $Sp(4)$ gauge theories [153, 356]. Their use in the study of connected diagram contributions to flavour-singlet meson correlation functions in the $Sp(4)$ theory coupled to $N_{\rm f}=2$ fundamental and $N_{\rm as}=3$ antisymmetric fermions dramatically improves noise control, and makes it possible to perform measurements in this sector of the theory [156]. Application to the same theory further provides a supporting tool for the spectral density approach to lattice spectroscopy [155].
@@ V.1 Masses and decay constants
We have evidence in our numerical results suggesting that in the limit of massless hyperquarks of the theory, confinement is taking place, with spontaneous breaking of the (approximate) non-Abelian global symmetry. [^8] As we have clear evidence of symmetry breaking, even in naive extrapolations to the massless limit, we do not discuss symmetric mass generation—see Ref. [370] and references therein. The resulting low-energy spectrum contains light pseudoscalar meson states identified as the pseudo Nambu-Goldstone bosons. Yet, despite the wide range of hyperquark masses we explored, our results suggest that we expect to see significant deviations from the predictions of leading order (LO) chiral perturbation theory, as none of the ensembles are close enough to the massless limit. We devote the next paragraphs to presenting this evidence and discussing its implications.
@@
In Fig. 10 we combine our measurements of the masses of ${\rm ps}$, ${\rm v}$, ${\rm s}$, ${\rm av}$, and ${\rm at}$ mesons, and present them in units of the decay constant of the ${\rm ps}$ meson, $f_{\rm ps}$. We display the results as a function of $am_{\rm PCAC}$. The figure highlights the contrast between the scaling properties of the ${\rm ps}$ meson, and all the other mesons. This observation further enforces the conclusion that the massless extrapolation of this theory is chirally broken, not IR conformal, and the ${\rm ps}$ mesons are indeed (light) Goldstone bosons. This therefore justifies the adoption of an analysis strategy that assumes confinement and spontaneous symmetry breaking of (approximate) global symmetry.
@@ V.3 More about vector mesons
Even more speculatively, this ratio, ${m_{{\rm v}}}/{f_{{\rm ps}}}$, can be used to provide a rough estimate of the on-shell coupling constant associated with the decay of the vector into two pseudoscalar mesons, $g_{\rm v\,ps\,ps}$, via the phenomenological KSRF relation, $g_{\rm v\,ps\,ps}=m_{\rm v}/\sqrt{2}f_{\rm ps}$ [367, 368]. In the continuum and massless limits, our numerical results yield $g_{\rm v\,ps\,ps}=4.656(80)$, which is smaller than the QCD value of $\sim 5.9$ [375] and the one in the $Sp(4)$ gauge theory with two fundamental fermions, $m_{\rm V}/\sqrt{2}f_{\rm PS}=5.72(18)(13)$ [144]. It would be interesting to perform a direct measurement of the coupling $g_{\rm v\,ps\,ps}$, which requires to study the theory with lower hyperquark masses.
@@ VI Summary and Outlook
Table 10: Ratio of the masses of the flavored mesons included in the continuum and massless extrapolations, to the pseudoscalar decay constant, $m_{M}/f_{\rm ps}$. The masses are measured by applying APE and Wuppertal smearing in the definition of the correlation functions. The mass of the first excited state of vector mesons, ${\rm v}^{\prime}$, is determined through a GEVP analysis that includes the two, inequivalent, vector and tensor meson operators. We also report the ratio of the masses of the first-excited and ground states of the vector meson, $m_{{\rm v}^{\prime}}/m_{\rm v}$.
@@ Appendix A More numerical results
In this appendix, we summarise all the new numerical results on lattice measurements of meson masses and decay constants which have been carried out for this work. We adopted two different strategies for the analysis of the correlation function. Measurement extracted from correlation functions defined with $Z_{2}\otimes Z_{2}$ stochastic wall sources and sinks, for all the ensembles listed in Table II.1, are tabulated in Tables  VI and  VI, and are discussed in Sect. V.1 and Appendix C. We also report the results for the PCAC (unrenormalised) hyperquark mass, $am_{\rm PCAC}$, and for the quantities, $m_{\rm ps}L$ and $f_{\rm ps}L$. To facilitate comparison with phenomenological models and applications to model building, we find it useful to express the measurements also in units of the decay constant of the pseudoscalar meson, $f_{\rm ps}$, in Table VI.
As explained in the body of the paper, we select a subset of the ensembles, which we use for the massless and continuum extrapolation, and apply a combination of APE and Wuppertal smearing, to repeat the mass measurements, performing also the extraction of the first-excited states of vector mesons with the GEVP method. The results are tabulated in Table VI, and discussed in Sects. V.2 and V.3. We express them in units of $f_{\rm ps}$, in Table VI, in which we also report the ratio between the first excited and ground states of vector meson.
@@ Appendix D PCAC mass
We conclude by repeating that, even if the dEFT fit to the data yields intriguing results, it remains a somewhat speculative exercise. Besides the aforementioned fact that we do not know whether a light scalar singlet is present in the spectrum, this analysis relies on the use of continuum relations, while we know that the values of $\beta$ available are such that significant corrections are expected to be for this analysis due to discretisation effects. Another concern might originate from the observation that the quality of the dEFT fit worsens as the $\beta$ value increases. Nevertheless, the results are encouraging, and it would be interesting to redo such analysis on lattice data that is closer to both the continuum and the massless limits.